IMPROVED EXPERIMENTAL LINE POSITIONS FOR THE (1,1) BAND OF THE b^{1}Σ^{+} - X^{3}Σ^{−} TRANSI-TION OF O_{2} BY INTRACAVITY LASER ABSORPTION SPECTROSCOPY

\underline{LEAH C. O’BRIEN}, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; EMILY C. O’BRIEN, JAMES J. O’BRIEN, Department of Chemistry & Biochemistry and Center for NanoScience, University of Missouri, St. Louis, MO 63121-4400.

We report improved experimental line positions for the (1,1) band of the b^{1}Σ^{+} - X^{3}Σ^{−} transition of O_{2}. Results are comparised with previous experimental measurements and predicted values. Additionally, a new method of producing vibrationally hot molecules for use in absorption spectroscopy of stable gas phase molecules is described.